OpenStudy (anonymous):

A solution of salt and water contains 70 grams of water per 140 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 35 milliliters of the solution. [Round off your answer to two decimal places]

ganeshie8 (ganeshie8):

\(\large \frac{70~grams}{140~millileters}\)

ganeshie8 (ganeshie8):

\(\large \frac{70~grams}{140~millileters} \times \frac{1~mole}{16~grams}\)

OpenStudy (anonymous):

N1V1=N2V2 N1= no. of moles before V1= volume before V2=volume after N2=no. of moles after now you can solve!!

OpenStudy (anonymous):

2. something

OpenStudy (ranga):

140ml of solution has 70 grams of water 35ml of solution will have how many grams of water? divide that by 16 grams to find the number of moles of water.

ganeshie8 (ganeshie8):

\(\large \frac{70~grams}{140~millileters} \times \frac{1~mole}{16~grams} \times 35 ~millileters\)

ganeshie8 (ganeshie8):

cancel watever u can, including units also

OpenStudy (anonymous):

2.1moles

OpenStudy (ranga):

check your calculations. should be 1.something

OpenStudy (ranga):

I have a question about one of the numbers in this problem: Shouldn't 1 mole of water weigh 18 grams instead of 16 grams as stated in the problem? Molecular weight of H20 = 2H + O = 2*1 + 16 = 18. Therefore, 1 mole of water weighs 18 grams.

ganeshie8 (ganeshie8):

ahh yes it should be 18g/mol^^

ganeshie8 (ganeshie8):

mistake in the question